OpenStudy (anonymous):

How would I rationalize 14/(3+sqrt2)

OpenStudy (anonymous):

\[\frac{14}{3+\sqrt2}\times \frac{3-\sqrt2}{3-\sqrt2}\] is a start

OpenStudy (whpalmer4):

\[\frac{14}{3+\sqrt{2}}\]multiply top and bottom by (3 - sqrt(2)) to give \[\frac{14}{(3+\sqrt{2})}*\frac{(3-\sqrt{2})}{(3-\sqrt{2})} = \frac{42-14\sqrt{2}}{9-2}\] and the rest of the steps should be obvious...

OpenStudy (anonymous):